OpenStudy (anonymous):

Choose one of the factors of 5x3 – 135 Answer All of the above 5 x2 + 3x + 9 x – 3

OpenStudy (anonymous):

x-3 is the answer!

OpenStudy (anonymous):

\[5x^3=135\] \[x^3=135/5\] \[x^3=27\] \[x=\sqrt[3]{27}\] \[x=3\] \[x-3=0\]

OpenStudy (anonymous):

no its not. i got it wrong

OpenStudy (anonymous):

try 5

OpenStudy (anonymous):

I don't know the right answer but (5) is WRONG.

OpenStudy (anonymous):

for people who have this in the future the answer is all the above

OpenStudy (anonymous):

is it ? @samiepeeler

OpenStudy (anonymous):

yes it is I just took it like 2 seconds ago @coltonbasham